solve by elimination method
4x-9y=14.5
7y-2x=-8.5

Answer by checkley77(12844)

4x-9y=14.5
7y-2x=-8.5 multiply by 2 & add.
-4x+14y=-17
4x-9y=14.5
----------------
5y=-2.5
y=-2.5/5
y=-.5 ans.
-----------
4x-9*-.5=14.5
4x+4.5=14.5
4x=14.5-4.5
4x=10
x=10/4
x=2.5 ans.